### v2

# Mesh-Learner: Texturing Mesh with Spherical Harmonics

Abstract
In this paper, we present a 3D reconstruction and rendering framework termed Mesh-Learner that is natively compatible with traditional rasterization pipelines. It integrates mesh and spherical harmonic (SH) texture (i.e., texture filled with SH coefficients) into the learning process to learn each mesh’s view-dependent radiance end-to-end. Images are rendered by interpolating surrounding SH Texels at each pixel’s sampling point using a novel interpolation method. Conversely, gradients from each pixel are back-propagated to the related SH Texels in SH textures. Mesh-Learner exploits graphic features of rasterization pipeline (texture sampling, deferred rendering) to render, which makes Mesh-Learner naturally compatible with tools (e.g., Blender) and tasks (e.g., 3D reconstruction, scene rendering, reinforcement learning for robotics) that are based on rasterization pipelines. Our system can train vast, unlimited scenes because we transfer only the SH textures within the frustum to the GPU for training. At other times, the SH textures are stored in CPU RAM, which results in moderate GPU memory usage. The rendering results on interpolation and extrapolation sequences in the Replica and FAST-LIVO2 datasets achieve state-of-the-art performance compared to existing state-of-the-art methods (e.g., 3D Gaussian Splatting and M2-Mapping). To benefit the society, the code will be available at https://github.com/hku-mars/Mesh-Learner.
I INTRODUCTION
Nowadays, 3D reconstruction and novel view synthesis are thriving research directions, as they can be applied in simulation task [1], robot’s navigation [2], entertainment [3] applications, and so on. In the past, traditional algorithms like multi-view stereo (MVS) [4] were commonly used in this field. Recently, researchers have turned to Neural Radiance Fields (NeRF) [5] to address the shortcomings of traditional 3D primitives such as point clouds and voxels, enabling the use of continuous implicit functions parameterized by neural networks. However, these methods suffer from the limitations of real-time rendering and extended training durations. Subsequently, 3D Gaussian Splatting (3DGS) [6] was proposed, offering significant improvements in rendering quality and training speed compared to NeRF. Existing works such as [7, 8, 9] have significantly improved the rendering quality and efficiency of NeRF and 3DGS compared to their original versions [6, 5]. However, they ignore a practical issue: inability to directly integrate with traditional rasterization pipelines. To visualize NeRF or 3DGS in rendering engines like Unity, users typically introduce third-party plugins, such as [10].
Figure 1: The rendering results of Mesh-learner on real-world FAST-LIVO2 dataset.
However, it suffers from artifacts [11] and fundamental rendering errors (e.g., occlusion error [12]). Furthermore, many applications already have extensive mesh models and require interaction with trained 3DGS or NeRF models. Therefore, there is a need to convert 3DGS or NeRF models into mesh format. However, these conversions not only introduce additional workload but also inevitably lead to losses in rendering quality and geometric integrity. Such limitations lead people to consider whether they can directly learn a texture from the existing mesh, so the learning outcome can be directly applied to a rasterization pipeline without conversion. This is based on the ability to obtain satisfactory meshes with the help of other lidar-based algorithms [13, 14].
To diminish mentioned gap between the algorithm outcome and the downstream tasks, we present Mesh-Learner, an end-to-end framework based on rasterization pipeline for both training and inference, obtaining realistic novel view images. Mesh-Learner assigns each mesh a SH texture, utilizes rasterization pipeline to determine which SH Textures are involved in rendering each frame and sample on them. At sample points, we propose a hybrid interpolation method between SH texels to obtain view-dependent colors. We extend the Elliptical Weighting Average (EWA) filter [15] from local texture space into world space on SH textures to mitigate aliasing artifacts in distant scenes. In addition, we introduce an adaptive SH density strategy to determine the optimal SH texture resolution each mesh during training, according to the varying complexity of details in training scenes.
Mesh-Learner is developed in LibTorch C++ [16] and CUDA C [17] with careful engineering optimization. Additionally, users of Mesh-Learner can seamlessly integrate the training outcomes (SH textures) into their rasterization pipeline for visualization, as Mesh-Learner’s inference (rendering) part can be implemented solely using OpenGL. In summary, the contributions of our work are:
  1. 1.
    A spherical-harmonic-augmented texture with a rasterization-pipeline-based training framework to enhance compatibility and capability of photorealistic rendering.
  2. 2.
    A hybrid interpolation method that leverages adjacent SH Texture’s information to represent view-dependent radiance.
  3. 3.
    A world space EWA filter to reduce aliasing for distant scenes.
  4. 4.
    An adaptive SH density adjustment strategy to find the optimal SH texture resolution of each mesh.
II RELATED WORKS
Neural Radiance Fields (NeRF) [5] have pioneered a new track in 3D reconstruction and Novel View Synthesis (NVS) by utilizing neural networks to represent a scene. It maps a 3D position and a view direction to RGB color and density, employing volumetric rendering to produce photorealistic novel view images. Plenoxels [18] attempts to model a scene without using neural networks; instead, it employs a sparse voxel grid to define a scene and place SH coefficients at each voxel. They utilize volume rendering to render. At each sample point, they calculate the color and opacity using trilinear interpolation of neighboring SHs, then integrate all sample points along the ray. Recently, 3DGS [6] utilizes numerous Gaussian kernels instead of a fixed voxel grid to represent scene geometry and also uses SHs on each Gaussian kernel to store view-dependent radiance. They use rasterization to splat Gaussian kernels onto the image plane, sort them based on depth, and blend all the Gaussians at each pixel to composite the color. In comparison to them, the rendering method in our work combines rasterization and interpolation, aiming to achieve compatibility with the traditional rasterization pipeline and enhance rendering efficiency.
Several works introduce textures into 3DGS. [19] assign each Gaussian a distinctive texture. During rendering, they trace a ray from the camera and calculate the intersection point of the hit Gaussians. Then, they query the RGBA texture of each intersected Gaussian and composite the color and alpha to obtain the final color. Another similar work is [20]. They introduce learnable texture maps on 2D Gaussian surfels, based on [21], enabling each primitive to enhance its ability to represent more complex and detailed appearances. They perform ray-surface intersection and calculate the texture coordinates for each surfel’s texture, then obtain the zeroth basis of Spherical Harmonics by bilinearly sampling the texture map. In summary, the above methods treat texture as a medium attached to Gaussians to store colors. Unlike them, we directly implement rendering and optimization on textures.
Few works integrate traditional rendering pipelines into the training and rendering process of NeRF or 3DGS. [22] first represented the NeRF model based on textured polygons. They exploit the rasterization pipeline with Z-buffer to achieve pixel-level parallelism, providing each pixel with neural features in texture form. Then, they implement a MLP in the fragment shader to map those features to colors. However, this method is not end-to-end. First, they need to train a NeRF with continuous opacity, then binarize it, and finally bake opacity and features into textures for rendering. In comparison, our method is end-to-end and prioritizes ’high compatibility with traditional pipelines’ as the primary objective. Once the training converges, the output (SH textures) can be imported with the corresponding meshes into a rasterization pipeline for rendering, eliminating the need for any intermediate processing.
III METHODOLOGY
Figure 2: The overview of our proposed system.
The overview of our proposed system is illustrated in Fig. 2. The goal is to attain each mesh’s SH texture capable of rendering realistic view-dependent radiance. The system inputs are scene mesh and sampled posed images. We obtain high-precision meshes by using existing lidar-based methods (e.g., ImMesh [13], FAST-LIVO2 [14] and M2-Mapping [23]). To initialize, we allocate a SH texture for each mesh whose resolution is calculated based on SH Density (in Sec. III-A). Then in rendering (forward), it is comprised of two parts:
  • Deferred rendering (in Sec. III-B): we rasterize meshes into screen space, save at each pixel’s sampling texture coordinates, view directions, and other information of meshes in the frustum into offscreen textures.
  • CUDA Renderer (in Sec. III-C): we read offscreen textures from the deferred rendering stage, then on each pixel’s sampling point on SH Textures, we implement hybrid interpolation (in Sec. III-D) between neighbor SH texels to get pixel colors. If a rendering view is distant, we instead utilize our proposed world space EWA filter (in Sec. III-E) on SH textures to obtain anti-aliasing results.
In backward, a CUDA kernel is employed to record which SH texels contribute to each output pixel, whether through hybrid interpolation or world space EWA. Subsequently, based on the recorded relationships, the gradient of each pixel is backpropagated to the SH texels that influence the color of that pixel. Throughout the training, we periodically employ an adaptive SH Density strategy (in Sec. III-F) for each mesh to determine the optimal SH Texture resolution for each.
III-A SH density
Enlightened by Plenoxels [18] that interpolates SHs in 3D grids, we interpolate SHs on 2D SH Textures. We regard SH Texture as a tangible entity attached to the mesh surface, thereby we can assign each texel center a specific world coordinate. A simple approach to rendering is to assign a uniform resolution config to all SH textures and implement bilinear interpolation at pixels’ sampling points. However, this can lead to suboptimal rendering quality, as it overlooks that each mesh varies in size and shape (obtuse, acute, or right triangles). Specifically, in areas with high-frequency details, smaller meshes may yield satisfied interpolation results. However, on larger meshes with the same texture resolution config, SH texels might be too sparse to accurately represent those high-frequency details.
Figure 3: Illustration of SH Texture’s resolution calculation and interpolation areas. Each mesh is allocated a dedicated SH texture whose resolution is calculated based on SH density and triangle’s shape and size. For example, we calculate a SH Texture’s resolution on an obtuse triangle, as in (a). We find the longest edge as the triangle base, and SH texture’s width/height is determined by dividing the triangle base/height length by the SH density. The SH texels whose centers lie outside the triangle are deemed invalid. (b): We define three regions for a SH texture to realize different interpolation strategies. Inside (green): valid area where any sample point can find four texels for bilinear interpolation; Corner (purple): valid area where sample points are above the highest valid SH texel center or on the lower-left side of the leftmost valid SH texel center, or on the lower-right side of the rightmost valid SH texel center. Edge (blue): the rest valid area.
To address this, each mesh has to be assigned with different texture resolutions according to their distinctive shape and size. We calculate them by using SH density. SH density is defined as: the distance between two horizontally or vertically adjacent SH texels in a texture, measured in meters. It can be defined in meters because each SH texel within the SH Texture can be assigned a 3D coordinate in world space. To determine texture resolutions, as shown in Fig. 3a, we first find the longest edge of a triangle mesh and consider it as the triangle base. Next, we calculate the height relative to this base. Given that we know the world space coordinates of the triangle vertices, we can compute the length of the triangle base and height in world space, expressed in meters. The arrangement of SH texture is aligned with its related mesh: the base of the SH Texture coincides with the base of the triangle, and the height of the SH Texture aligns with the direction of the triangle’s height. To calculate the horizontal resolution of the texture, we divide the triangle’s base length by the SH density to determine the number of texels that can be arranged along the triangle base, in other words, the resolution of the SH texture in the horizontal direction. Similarly, we divide the triangle’s height by the SH density to obtain the number of SH texels that can be accommodated in the vertical direction, corresponding to the texture’s vertical resolution. It is evident that, since the texture is rectangular when aligned with a triangle, there will inevitably be texels whose centers lie outside of the triangle boundary. We mark these texels as invalid, and they do not participate in subsequent training.
Figure 4: Illustration of world space EWA and the calculation of world space coordinates of SH texels. (a): In the Inside region (red circle), we evaluate texels in the current texture; in the Corner region (blue circle), we evaluate texels in neighbor textures that share the same triangle vertex; in the Edge region (green circle), we evaluate texels in nearby textures that share the same edge. (b): The calculation of world space coordinates for SH texels.
Next, we calculate the world coordinates of each valid SH texel in SH Textures. As illustrated in Fig. 4b, red dots represent valid SH texel centers, while blue lines indicate the calculation steps. We sequentially traverse and compute the world coordinates of all valid SH Texels in the bottom row, moving to the row above and continuing this process until we reach the top row. This calculation is feasible because we know the world coordinates of the triangle vertices. We also know the direction and distance from the triangle vertices to the SH Texels in world space, as well as the distances between every two adjacent SH Texels (i.e. SH Density). We begin at the bottom-left corner. First, we add the initial step vector $\mathbf{V}_{0}$ on the bottom-left triangle vertex ${}^{W}\!\mathbf{p}_{a}$ to obtain the world coordinate of the bottom-left-most SH Texel $T_{0}$; then by adding the horizontal update vector $\mathbf{V}_{h}$ to $T_{0}$, we can find the world coordinates of its right-hand-side adjacent SH Texel $T_{1}$. Similarly, we apply the vertical update vector $\mathbf{V}_{v}$ on $T_{1}$ to get the world coordinate of $T_{2}$. We denote the triangle vertices as ${}^{W}\!\mathbf{p}_{a}$, ${}^{W}\!\mathbf{p}_{b}$, ${}^{W}\!\mathbf{p}_{c}$, and ${}^{W}\!\mathbf{p}_{d}$ is the triangle foot. The distance between two vertically or horizontally adjacent texels is the SH density, denoted as $\mathcal{D}$. So the horizontal $\mathbf{V}_{h}$ and vertical update vector $\mathbf{V}_{v}$ are calculated as follows:
$\mathbf{V}_{h}=\frac{\mathbf{{}^{W}\mathbf{p}_{b}-{}^{W}\mathbf{p}_{a}}}{\|\mathbf{{}^{W}\mathbf{p}_{b}-{}^{W}\mathbf{p}_{a}}\|}\mathcal{D},~{}~{}\mathbf{V}_{v}=\frac{\mathbf{{}^{W}\mathbf{p}_{c}-{}^{W}\mathbf{p}_{d}}}{\|\mathbf{{}^{W}\mathbf{p}_{c}-{}^{W}\mathbf{p}_{d}}\|}\mathcal{D}$
Lastly, the initial step vector is: $\mathbf{V}_{0}={}^{W}\!\mathbf{p}_{a}+\frac{\mathbf{V}_{h}}{2}+\frac{\mathbf{V}_{v}}{2}$. We repeatedly add updated vectors to get all valid SH Texels’ world coordinates.
III-B Deferred rendering
After initialization, we first perform deferred rendering at the beginning of each training step. This stage is not for rendering color, but for acquiring meshes in the current frustum and the information related to pixels and SH texels. The inputs for this stage include vertices, camera poses, and dummy textures. The outputs consist of sampling texture coordinates, EWA’s ellipse coefficients, level of detail (LOD) level, view direction, and the world space coordinates at each pixel’s sampling point (to be illustrated later). To obtain sampling texture coordinates of each pixel, we do not send all actual SH textures of all meshes to the GPU; instead, we use dummy textures. This is based on the facts that: 1) The combined memory footprint of all textures in the scene is too large to fit into GPU memory. 2) We only need texture coordinates at this stage, not sampling on textures. 3) Many textures share the same resolution configuration. 4) The number of distinct resolution configurations is orders of magnitude smaller than the total number of SH textures. To obtain the view directions and the world space coordinates of the sampling points, we employ barycentric interpolation during rasterization to interpolate the corresponding attributes of the triangle vertices. To get EWA’s ellipse coefficients, we retrieve screen space pixel gradients and compute the ellipse coefficient in shader as outlined in [15]. We also obtain the LOD level of each pixel in shader. All results are passed OpenGL’s depth test and saved to offscreen GPU textures.
III-C CUDA Renderer
We use CUDA-OpenGL interoperation to directly access the offscreen textures created by the deferred rendering stage on the GPU side. Based on that, for each pixel, we can identify which SH Texture is related to and the associated sampling texture coordinates. So, we can identify which region of the SH Texture this pixel belongs to (Sec. III-D). Following this, we determine whether to use our hybrid interpolation or world space EWA filter to render based on each pixel’s LOD level. For example, if the LOD level is greater than 1, we activate world space EWA; otherwise, we use hybrid interpolation.
III-D Hybrid SH Interpolation
We have texture coordinates at each pixel’s location, but relying only on bilinear interpolation on SH texels is not enough, especially when sampling at texture edges and corners. The bilinear method may not suffice as it requires four texels, which might not be available in this area. So it’s natural to attend to adjacent SH textures.
First, we define the Valid Area: as illustrated in Sec. III-A, a valid SH texel lies within the scope of the triangle mesh. Similarly, a valid area is where sampling points are located within the triangle mesh’s scope. Then, we divide a SH Texture into three regions: Inside, Corner, and Edge (as shown in Fig. 3b). Inside: a valid area where we find four valid SH Texels surrounding a sample point to conduct bilinear interpolation; Corner: a valid area that is above the highest valid SH Texel center, or located on the lower-left side of the left-bottom-most valid SH texel center, or the lower-right side of the right-bottom-most valid SH texel center; Edge: the remaining valid area.
We implement different interpolation strategies in different regions. For a sampling point in the Inside region, since we can find nearby four texels, we use bilinear interpolation. If it’s in the Edge region, we cannot find four texels in self SH Texture for bilinear interpolation. The possible number to find can be one, two, or three, depending on the location of the sampled texture coordinate within this region. We then locate one additional nearest SH texel in the neighbour SH texture who shares the same edge and use Inverse Distance Weighting (IDW) to interpolate among selected texels:
$u(\mathbf{x})=\begin{cases}\frac{\sum_{i=1}^{N}w_{i}(\mathbf{x})u_{i}}{\sum_{i=1}^{N}w_{i}(\mathbf{x})},&\text{ if }d\left(\mathbf{x},\mathbf{x}_{i}\right)\neq 0\text{ for all }i\\ u_{i},&\text{ if }d\left(\mathbf{x},\mathbf{x}_{i}\right)=0\text{ for some }i\end{cases}$ (1)
where
$w_{i}(\mathbf{x})=\frac{1}{d\left(\mathbf{x},\mathbf{x}_{i}\right)^{0.9}}$
where $\mathbf{x}$ represents the sampled location, $\mathbf{x}_{i}$ denotes the locations of selected SH Texels, and $\mathbf{u}_{i}$ indicates the corresponding SH coefficients. The distance function $d(\mathbf{x},\mathbf{x}_{i})$ represents the Euclidean distance in world space, $N$ indicates the number of SH Texels to be interpolated. If the sampled texture coordinate is in the Corner region, plus the neighboring SH Texels in self SH texture, we locate one nearest SH texel in each neighboring SH texture and apply IDW to interpolate among these texels. We can use IDW because each SH Texel has been assigned a world space coordinate during initialization, and the world space coordinate of the texture sample point (i.e., interpolation center of IDW) can be obtained from deferred rendering, as we mentioned in (Sec. III-B). We choose IDW because the computational complexity is relatively low, which supports a large amount of pixel calculations while yielding acceptable interpolation results.
III-E World Space EWA Filtering
The results from the aforementioned method are quite commendable when cameras capture close-up scenery. However, a noticeable aliasing artifact remains when cameras face distant landscapes. To mitigate this, people often use MipMap [24] or other filtering methods. Considering the variability in the shape and size of meshes, the diverse camera viewpoints, and aiming to leverage SH Texels on neighboring meshes, we propose using anisotropic filtering in world space. We extend EWA filtering [15] from local texture space to world space. This allows us to evaluate whether SH Texels in the current SH texture and adjacent SH textures are within the ellipse range. We then use in-range SH Texels to calculate the elliptical weighted average.
The EWA’s ellipse coefficients are calculated during the deferred rendering stage; they are originally in local texture space and can be transformed into world space. This is based on the following: 1) we have world space coordinates of the sampled point (i.e., the ellipse center), obtained during the deferred rendering stage. 2) we can calculate the ellipse’s axis lengths in world space by multiplying the texel distance by the SH density. To determine whether the SH texels on neighbouring or self-textures lie within the ellipse in world space, we compute the world coordinates of the SH texels, which were determined in Sec. III-A.
So far, since both the ellipse equation and SH texels’s position are in the same world space, we can determine if a SH texel on the current SH texture or its neighbouring SH textures falls within the ellipse’s range by substituting the coordinates of the SH texel into the ellipse equation. Note that this algorithm assumes that the current mesh and adjacent meshes lie on the same plane. However, this assumption may not always be valid. Therefore, we compare the normals of the current and adjacent meshes; if the angle between the two normals exceeds $15$ degrees, we skip EWA for that adjacent mesh. The strategy for selecting neighbouring texels is similar to that of the hybrid interpolation method: when the SH texture sample point is in the Inside region, we apply EWA solely to the current texture; if in the Edge region, we implement EWA on both the current texture and the neighbouring texture that shares the same edge; and if in the Corner region, EWA is applied to the current texture and the neighbouring textures that share the same vertex.
III-F Adaptive SH Density Strategy
Although the resolutions of SH textures differ due to varying mesh sizes and shapes, their SH densities remain the same. It is not ideal to set SH density globally constant, as the complexity of scene elements is inconsistent across scene regions. If SH density is too low, high-frequency details cannot be adequately represented through interpolation between sparse SHs. Conversely, setting SH density too high results in aliasing, as it results in the SH Density, particularly in distant scenes, exceeding half the camera sampling rate (according to the Nyquist-Shannon Sampling Theorem [25]), leading to aliasing artifacts.
Notation: Initial SH Density $\rho_{0}$, number of converged mesh $N_{cvg}$, density update step $d$, threshold of max patiance number $T_{patiance}$, total number of meshes $N_{total}$.
Input: -- Current round PSNR average ${\mathcal{A}_{i}}$, PSNR variance ${\mathcal{V}_{i}}$, and retry times $C_{i}$;
-- Last round PSNR average ${{\bar{\mathcal{A}}_{i}}}$, PSNR variance ${{\bar{\mathcal{V}}_{i}}}$, and SH Density ${\bar{\rho_{i}}}$.
-- Mesh set $\mathcal{M}_{i}$ and termination percentage of converged mesh $\epsilon_{T}$.
Output: new SH Densities $\rho_{i}$.
1
[…]
3 while $\frac{N_{cvg}}{N_{total}}<\epsilon_{T}$ do
4       if $\mathcal{M}_{i}$ is Converged then
5            continue
6       end if
7       if $\mathcal{M}_{i}$ is first time update density then
8             $\rho_{i}:=\bar{\rho}_{i}+d$
9             Recalculate SH Texture’s resolution
10       end if
11
12      if ${\mathcal{A}_{i}}>{{\bar{\mathcal{A}}_{i}}}$ OR ${\mathcal{V}_{i}}<{{\bar{\mathcal{V}}_{i}}}$ then
13             $\rho_{i}:=\bar{\rho}_{i}+d$
14             Recalculate SH Texture’s resolution
15       end if
16      else
17             if $C_{i}<T_{patiance}$ then
18                   $C_{i}=\bar{C}_{i}+1$ continue
19             end if
20            $\rho_{i}:=\bar{\rho}_{i}$
21             $N_{cvg}=N_{cvg}+1$
22             $\mathcal{M}_{i}$ Set Converged
23       end if
24
25 end while
Algorithm 1 Adaptive SH Density Strategy
Inspired by [9], we propose an adaptive strategy to address this issue: we assign each mesh a globally low initial SH density and adjust the SH Density changing direction (i.e., increase its density) through a metric method, aiming to adjust each SH density for improved metrics until no further enhancement is achievable. Specifically, after the initial SH density assignment, we train the first round to converge and then evaluate the averages and variances of PSNR based on all input views for each mesh separately. Subsequently, we update SH Densities according to each mesh’s initial density update direction, recalculate SH Texture resolutions, and begin a second round of training. Once it converges, we can obtain another set of metrics (i.e., average and variance of PSNR) of each mesh. We compare these metrics with the previous counterparts; if the average PSNR of a mesh increases or the variance of a mesh’s PSNR decreases, we consider the current update direction to be correct. Otherwise, we consider there’s no need for further increasing the SH Density. Considering the stochastic nature of the training process, we introduce a ”patience” parameter. If the metric does not improve, we allow the mesh another opportunity to optimize in the next training round. Should a mesh’s metrics still not improve under these circumstances, we consider it converged, as it can no longer be optimized to enhance its metrics. In Algorithm 1, $\mathcal{M}_{i}$ refers to a single mesh within the set $\mathcal{M}$ (i.e., all the meshes in a scene), and similarly, $C_{i},\mathcal{A}_{i},\mathcal{V}_{i},\bar{\mathcal{A}}_{i},\bar{\mathcal{V}}_{i},\rho_{i},\bar{\rho_{i}}$ refers to an attribute of a single mesh.
III-G Optimization
To recover scene radiance, we use smooth $L1$ Loss ${\mathcal{S}}_{L_{1}}(\cdot)$ [26] between the ground-truth pixel color $\boldsymbol{C}_{i}$ and rendered color $\hat{\boldsymbol{C}}_{i}$, where $i$ is the index of pixels:
${Loss}=\sum_{i}{\mathcal{S}}_{L_{1}}(\boldsymbol{C}_{i}-\hat{\boldsymbol{C}}_{i}),$ (2)
where:
${\mathcal{S}}_{L_{1}}(x)=\begin{cases}0.5x^{2}&\text{if }|x|<1\\ |x|-0.5&\text{otherwise}\end{cases}$ (3)
IV Experiment
In this section, we test and compare the rendering outcome of our framework on interpolation and extrapolation sequences under synthetic (Replica [27]) and real dataset (Fast-LIVO2 Dataset [14]).
IV-A Experiments Settings
IV-A1 Baselines
Our work is compared to other state-of-the-art methods. For NeRF-based methods, we include InstantNGP [28] and M2-Mapping [23]. For 3DGS-based methods, we select 3D Gaussian Splatting (appearance-oriented) [6] and MonoGS (RGBD-based) [29].
IV-A2 Metrics
To evaluate rendering quality, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to compare the rendering results with the corresponding ground truth image.
IV-B Replica Dataset
The Replica dataset [27] is a synthetic dataset of indoor scenes. It generates images from specified viewpoints by rendering scene models using OpenGL. We employ the extrapolation and interpolation trajectories defined by M2-Mapping [23] for testing. As shown in Table II and Figure 8, our proposed framework captures more precise details in both remote and nearby scenes, for both extrapolation and interpolation trajectories.
TABLE I: Quantitative results on the FAST-LIVO2 dataset.
Metrics Methods Station SYSU CBD Avg.
[…]
3DGS 26.859 28.798 22.078 25.825
Ours 25.453 26.010 25.409 25.624
[…]
3DGS 0.853 0.825 0.789 0.822
Ours 0.833 0.798 0.858 0.829
[…]
3DGS 0.142 0.279 0.308 0.243
Ours 0.172 0.286 0.201 0.219
Figure 5: Comparison results between using and not using our proposed world space EWA filter. The noise here comes from an insufficient sampling rate, which results in an inadequate number of SH texels involved in the interpolation, causing the resulting colors to diverge from the GT.
Figure 6: Comparison of our Adaptive SH Density Strategy.
Figure 7: The qualitative results on FAST-LIVO2 Datasets.
Figure 8: Comparison results on Replica Dataset’s Office-0. The results demonstrate that our work captures more details in both extrapolation and interpolation sequences. In the first row, there’s a zoomed-in view from an extrapolation perspective of a distant clock in the scene. In our results, the small numbers are clear enough to read, while the outcomes of other works are unreadable. In the second row’s zoomed-in view from an extrapolation perspective, our result has sharper edges (blue pad) and more detail (white books). From an interpolation view in the last row, the pattern on the wall in our result is clearer and sharper compared to 3DGS and M2-Mapping.
TABLE II: Quantitative interpolation(I) and extrapolation(E) rendering results on the Replica Dataset.
Metrics Methods Office-0 Office-1 Office-2 Office-3 Office-4 Room-0 Room-1 Room-2 Avg.
SSIM(I)$\uparrow$ InstantNGP 0.981 0.980 0.963 0.960 0.966 0.964 0.964 0.967 0.968
3DGS 0.987 0.980 0.980 0.976 0.980 0.977 0.982 0.980 0.980
MonoGS 0.985 0.982 0.973 0.970 0.976 0.971 0.976 0.977 0.975
M2-Mapping 0.985 0.983 0.973 0.970 0.977 0.968 0.975 0.977 0.976
Ours 0.990 0.986 0.980 0.973 0.976 0.976 0.988 0.982 0.981
PSNR(I)$\uparrow$ InstantNGP 43.538 44.340 38.273 37.603 39.772 37.926 38.859 39.568 39.984
3DGS 43.932 43.237 39.182 38.564 41.366 39.081 41.288 41.431 41.010
MonoGS 43.648 43.690 37.695 37.539 40.224 37.779 39.563 40.134 40.034
M2-Mapping 44.369 44.935 39.652 38.874 41.318 38.541 40.775 40.705 41.146
Ours 44.689 43.980 41.132 38.723 41.352 39.267 42.876 41.784 41.725
LPIPS(I)$\downarrow$ InstantNGP 0.046 0.069 0.096 0.087 0.089 0.079 0.094 0.094 0.082
3DGS 0.040 0.075 0.069 0.068 0.065 0.060 0.057 0.075 0.064
MonoGS 0.031 0.048 0.061 0.057 0.052 0.055 0.050 0.056 0.051
M2-Mapping 0.031 0.044 0.055 0.050 0.050 0.062 0.056 0.053 0.050
Ours 0.025 0.028 0.027 0.032 0.036 0.022 0.023 0.028 0.028
SSIM(E)$\uparrow$ InstantNGP 0.972 0.961 0.934 0.938 0.952 0.918 0.936 0.941 0.944
3DGS 0.936 0.897 0.924 0.917 0.925 0.881 0.915 0.919 0.914
MonoGS 0.974 0.961 0.945 0.942 0.950 0.912 0.942 0.946 0.947
M2-Mapping 0.980 0.976 0.960 0.964 0.970 0.955 0.963 0.965 0.967
Ours 0.987 0.979 0.973 0.971 0.978 0.964 0.975 0.970 0.974
PSNR(E)$\uparrow$ InstantNGP 39.874 39.120 31.274 32.135 34.458 32.587 33.024 32.266 34.341
3DGS 31.220 29.959 27.411 26.442 28.324 27.541 28.429 27.139 28.307
MonoGS 39.197 38.818 29.740 29.664 31.632 29.949 31.126 30.621 32.593
M2-Mapping 41.965 42.215 35.056 37.465 38.667 36.427 37.294 36.722 38.226
Ours 42.480 41.751 35.202 36.154 38.821 35.510 36.872 36.531 37.915
LPIPS(E)$\downarrow$ InstantNGP 0.085 0.117 0.176 0.185 0.149 0.180 0.162 0.156 0.151
3DGS 0.117 0.177 0.167 0.181 0.155 0.204 0.170 0.174 0.168
MonoGS 0.061 0.104 0.136 0.137 0.108 0.170 0.123 0.117 0.120
M2-Mapping 0.046 0.062 0.084 0.081 0.066 0.114 0.084 0.087 0.078
Ours 0.027 0.058 0.061 0.066 0.051 0.053 0.046 0.061 0.047
IV-C FAST-LIVO2 Dataset
For real-world scenarios, we evaluate our framework on the FAST-LIVO2 datasets [14]. The localization results from FAST-LIVO2 are used as ground truth poses. To evaluate our algorithm’s interpolation results, we split the pose-image-pair sequence generated by FAST-LIVO2 into the train and test set. Specifically, for a sequence, we extract every 8th pose-image-pair for the test set. The remains are for the train set. The quantitative results are shown in Tab. I. Our work shows promising rendering results compared to 3DGS [6] and surpass the state-of-the-art in certain large-scale scenes. The qualitative results on the interpolation sequence are in Fig. 7. Our world space EWA filter can mitigate aliasing artifacts in distant view, in Fig. 5. What’s more, we can render high-frequency regions well with high SH density while maintaining low SH density in mono color area, as shown in Fig. 6.
IV-D Results
Rendering with hybrid interpolation alone on SH Textures under fixed SH density does not yield satisfactory results due to aliasing artifacts caused by insufficient sampling rates. However, adopting the World Space EWA Filter and Adaptive SH Density Strategy can significantly improve the results. As shown in Fig. 5, in the distant regions of the image, our world space EWA filter averages SH texels within the ellipse at each pixel’s SH texture sampling point, resulting in smoother outcomes with less noise. We also implement the Adaptive SH Density Strategy to further enhance our rendering quality, as illustrated in Fig. 6. Without it, rendering scenes with complex textures requires setting the global SH Density very high. This leads to aliasing artifacts in distant regions where the SH Density surpasses the camera’s sampling rate (purple box). Similar artifacts, such as Moiré patterns, can also be observed within the red box. After implementing our adaptive strategy, we achieve a higher density in areas with complex textures (red box), which leads to improved rendering quality while keeping a lower SH Density in areas with simple textures, thereby eliminating aliasing artifacts.
V Conclusion and Future Works
V-A Conclusions
In this work, we propose a 3D reconstruction and rendering framework called Mesh-Learner, which is designed to learn view-dependent radiance while ensuring high compatibility with existing rasterization pipelines. We assign each mesh a SH Texture whose resolution is calculated by SH density, and employ hybrid interpolation on SH Textures at the sampling point of each pixel to obtain view-dependent color. We propose a World Space EWA Filter to reduce artifacts caused by insufficient sampling rates in remote scenes. Additionally, we implement an Adaptive SH Density strategy to determine the optimal SH Density for each SH Texture, further enhancing rendering quality while minimizing aliasing. Our work is evaluated on both synthetic (i.e., Replica [27]) and real datasets (i.e., FAST-LIVO2 [14]), validating the effectiveness of our methods and demonstrating their potential for downstream tasks in robotics.
V-B Limitations and future works
One primary limitation of our work is that we critically depend on the quality of the mesh. Specifically, in a scene with occlusion relationships, if the mesh of the foreground object is incomplete, it can lead to the colors that should belong to the foreground object being erroneously optimized onto the mesh of the background object in the training images. As a result, the incorrectly learned colors of the background object may be rendered in other viewpoints, producing erroneous outcomes. To address this, our future work will focus on optimizing mesh vertices with SH Textures to refine defective meshes and achieve better reconstruction and rendering results.